Prove: There is no positive integer n such that n^2+n^3=100.



1
Expert's answer
2021-09-24T14:51:38-0400

Both functions "f1(n)=n^2" and "f2(n)=n^3" is increasing if n is positive integer, which means their sum is increasing too. 5^3=125, so, the easiest way to prove the statement is to calculate "f(n)=n^2+n^3" for integers from 1 to 4.

n=1: "f(1) = 2"

n=2:"f(2) = 12"

n=3:"f(3) = 36"

n=4:"f(4) = 80"

If n is negative integer, then n^2 + n^3 ≤ 0, so it cannot be equal to 100.

If n = 0 then f(0) = 0.

The statement has been proven.